The force F experienced by the charge Q=2c moving with the velocity V with a magnitude B is given by F=qv×B





Given





V= 2i+4j+6k





F=4i-20j+12k





What then is B





If Bx=By

1
Expert's answer
2022-01-23T16:13:57-0500

Given:

"{\bf v}= 2\hat i+4\hat j+6\hat k"

"{\bf F}= 4\hat i-20\hat j+12\hat k"

"q=2e"

"B_x=B_y=B"


The Lorentz force

"{\bf F}=q{\bf v}\times {\bf B}=2e\begin{vmatrix}
 \hat i & \hat j &\hat k \\
 2 & 4 &6\\
B& B& B_z
\end{vmatrix}"

"{\bf F}=2e[(4B_z-6B)\hat i-(2B_z-6B)\hat j-2B\hat k]"

Comparing with given force, one has

"B=\frac{12}{-4e}=\frac{3}{|e|}"

"B_z=\frac{10}{e}+\frac{3B}{e}=-\frac{19}{|e|}"

Finally,

"{\bf B}=\frac{3}{|e|}(1,1,-19/3)"